# Agentic SLMs: Hunting Down Test Smells

### v1

@@ 4.3 Feedback Loop
The feedback loop was triggered for detection primarily in cases where Agents 2 and 4 disagreed with Agents 1 and 3, respectively (Figure 1). For detecting test smells, Gemma 2 9B and Phi 4 14B utilized the feedback loop in 1 and 9 test smell instances, respectively. In 2 instances, Phi 4 14B was able to correctly identify a test smell due to the feedback loop. Llama 3.2 3B did not require the feedback loop for detecting any test smells. DeepSeek-R1 14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop. We employed a three-iteration feedback loop, which was sufficient for Agents 1 and 2 to reach a consensus on detecting test smells. For Agents 3 and 4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations. By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
@@ 4.4 Pull Requests
To assess the acceptance of the refactorings generated by Phi 4 14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects. As of this writing, five of these pull requests have been accepted and integrated: two in janusgraph (addressing the Assertion Roulette and Exception Handling test smells), and one each in opengrok (Duplicate Assert), jenkins (Duplicate Assert), and lettuce (Magic Number Test). One pull request was rejected in data-transfer-project because contributors noted that the additional message in each assertion did not improve their code. Four pull requests remain open: one in janusgraph and three in testcontainers-java, all related to Duplicate Assert and Magic Number Test smells.
4.5 Proprietary and Larger Models
In March 2025, we compared Phi 4 14B with four agents and pass@5 against proprietary models—o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental—each using a single-agent setup. o1-mini achieved a refactoring performance of 66.7%, which was inferior to Phi 4 14B with four agents. Evaluating o3-mini-high on 25 test smells that o1-mini failed to refactor yielded a 40% success rate. To further assess difficult cases, we tested Gemini 2.5 Pro Experimental, the top-ranked model on LLM Arena at the time of writing, which successfully refactored 8 out of 37 test smells that Phi 4 14B (with four agents and five attempts) failed to fix, representing a 5.3% improvement. These results indicate that Phi 4 14B, when paired with agentic workflows, can rival proprietary state-of-the-art models while remaining cost-effective and executable on consumer-grade hardware.
@@ 4.6 Prompts
We refined the prompts iteratively. For instance, in the definition of Assertion Roulette, the phrase “has more than one assertion” proved to be more precise and effective than the commonly used “has multiple non-documented assertion” found in the literature [3]. Early experiments overlooked prompt minimization, resulting in suboptimal outcomes – consistent with Hsieh et al. [21]. Providing Agent 4 with the original and refactored code plus Agent 3’s full explanation overloaded the model and reduced its effectiveness. Similarly, overly detailed smell definitions negatively impacted performance. Although not formally tested, we observed that large prompts hindered reasoning. Streamlining definitions and instructing Agent 4 to respond concisely improved efficiency. These findings align with prior studies showing that excessive or irrelevant context degrades LLM performance [22]. In some cases, we found that sentence order influenced model reasoning, supporting observations from previous work [23].
@@ 4.7 Other Languages
We evaluate whether the same setup could detect and refactor test smells in other programming languages. Using Phi 4 14B with a single agent, we analyzed the popular python-oauth2 GitHub project. The model successfully removed over 10 instances of the Assertion Roulette, Exception Handling, and Magic Number test smells across multiple test cases. Furthermore, we evaluated Phi 4 14B on the testify GitHub project using Golang, demonstrating its effectiveness in eliminating the Assertion Roulette test smell. Using a single agent, the model successfully refactored the code by adding explanatory messages to each assertion. These preliminary results indicate that the proposed approach is promising for detecting and refactoring test smells across various programming languages.
@@ 4.8 Larger Test Cases
As previously noted, 89% of the test cases in the 11 open-source projects contain at most 30 lines of code (LOC). To further evaluate Phi 4 14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects. We employed the same four-agent setup to detect and refactor these test smells. Although the strategy successfully identified 89% of the test smells, it showed a decline in performance compared to previous results. The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
@@ 4.9 Threats to Validity
One potential threat to internal validity is data leakage [24], where test smells analyzed in this study may be part of the foundation models’ training data. To mitigate this, we applied Metamorphic Testing. We limit our analysis to test code with a maximum of 30 lines of code (LOC). While this constraint excludes longer test cases, a substantial number of test cases still fall within this criterion, ensuring a representative evaluation. Another factor is prompt design, which can lead to generic responses instead of targeted explanations. To minimize variability, we used uniform and concise prompts across all evaluations, which were carefully reviewed by three authors of this paper. Validating alignment between the model output and test smell definitions is also challenging. To ensure accuracy, three authors independently reviewed selected responses, verifying whether the refactored code adhered to the intended definitions. Construct validity is limited by the small number of examples for each test smell, though they are real cases from GitHub repositories.
@@ 5 Related Work
Aljedaani et al. [5] compiled a comprehensive catalog of 22 test smell detection tools, the majority of which are designed for Java, SmallTalk, C++, and Scala, and 4 refactoring tools. Soares et al. [12] conducted a mixed-methods analysis involving 485 Java projects, exploring the adoption of JUnit 5 features to improve test code quality. Wang et al. [25] propose a tool called PyNose to detect 18 types of test smells in Python. Virgínio et al. [26] propose a tool to detect test smells in Java. Lambiase et al. [27] introduce DARTS, an IntelliJ plugin that detects and refactors three test smells: Eager Test, General Fixture, and Lack of Cohesion of Test Methods. Pontillo et al. [6] proposed a machine learning (ML)-based approach to detect test smells, focusing on four specific types. Lucas et al. [28] investigated the capability of three Large Language Models to detect test smells across multiple programming languages.
We investigate the effectiveness of agentic workflows for detecting and refactoring test smells in Java test cases. Using LangChain and Ollama, our approach coordinates up to four agents to automate these tasks. Results show that Phi 4 14B, an open model, achieves strong refactoring performance, outperforming traditional single-prompt methods in robustness and flexibility. Notably, it runs efficiently on a MacBook Pro M3 with 18GB of RAM, showing that such workflows are viable on consumer-grade hardware and perform comparably to proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental. A key advantage of our approach is its extensibility: new smells can be added via natural language definitions. Preliminary experiments also confirm successful application to Python and Golang, highlighting its cross-language potential. Unlike conventional methods that only detect smells or suggest code changes, our framework supports interactive conversations and explanatory feedback, helping developers understand the rationale behind each refactoring and increasing trust in automated suggestions.
@@ 6 Conclusions
In this study, we evaluated small, open language models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents. The evaluation covered 150 instances of 5 common test smell types from real-world projects. In addition to detection, we assessed each model’s ability to perform automated refactorings. Phi 4 14B, Gemma 2 9B, and Llama 3.2 3B successfully detected nearly all test smell instances, achieving a pass@5 of 96% with four-agent setups. Phi 4 14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. To assess real-world applicability, we submitted 10 pull requests with code refactored by Phi 4 14B. Five were merged, one was rejected, and four remain open – demonstrating the practical utility of our approach in live development environments.
Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort. However, for specific smells such as Assertion Roulette, a single agent performed better, indicating that workflow design should be tailored to the smell type. For practitioners, agentic workflows provide an automated and practical solution to maintain test code quality. For researchers, this study highlights the promise of integrating foundation models into software engineering tasks, offering a scalable path for automating test smell detection and refactoring in diverse programming environments.

### v2

@@ 4.4 Feedback Loop
The feedback loop was triggered when Agents2,4 disagreed with Agents1,3, respectively (see Figure 1). For detecting test smells, Gemma-2-9B and Phi-4-14B utilized the feedback loop in 1 and 9 test smell instances, respectively. In 2 instances, Phi-4-14B was able to correctly identify a test smell due to the feedback loop. Llama-3.2-3B did not require the feedback loop for detecting any test smells. DeepSeek-R1-14B successfully detected 3 instances of the Duplicate Assert test smell following a feedback loop. We employed a three-iteration feedback loop, which was sufficient for Agents1,2 to reach a consensus on detecting the majority of test smells. For Agents3,4 to reach a consensus on refactoring test smells, only 7.3% of the cases required more than three iterations. By allowing additional iterations, we achieved consensus in all cases by the eighth iteration, except for a single instance of the Conditional Test Logic test smell.
@@ 4.5 Pull Requests
To assess the acceptance of the refactorings generated by Phi-4-14B, which achieved the best performance in our study, we submitted 10 pull requests, ensuring at least one for each test smell type across five different open-source projects. As of this writing, six of these pull requests have been reviewed and merged by human maintainers, who deemed them valuable: two in janusgraph (Assertion Roulette and Exception Handling), two in lettuce (Magic Number and Duplicate Assert), and one each in opengrok and jenkins (both addressing Duplicate Assert). Two pull request were rejected because contributors noted that the additional message in each assertion did not improve their code. Two pull requests remain open.
4.6 LLMs
In July 2025, we also run o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent setup on 150 test smells. Claude-4-Sonnet, o3, Gemini-2.5-Pro achieved a refactoring performance (pass@1) of 59.3%, 74.7%, and 72.0%, respectively—though at significantly higher computational cost.
We allowed Phi-4-14B with four-agents to make up to five attempts, rather than limiting it to just one. With pass@5 [9], it successfully detected 96% and refactored 75.3% of the test smell instances. These results indicate that the open Phi-4-14B, when paired with agentic workflows, can rival proprietary state-of-the-art LLMs while remaining cost-effective and executable on consumer-grade hardware.
@@ 4.7 Prompts
We refined the prompts iteratively. For instance, in the definition of Assertion Roulette, the phrase “has more than one assertion” proved to be more precise and effective than the commonly used “has multiple non-documented assertion” found in the literature [2]. Early experiments overlooked prompt minimization, resulting in suboptimal outcomes – consistent with Hsieh et al. [10]. Providing Agent4 with the original and refactored code plus Agent3’s full explanation overloaded the model and reduced its effectiveness. Similarly, overly detailed smell definitions negatively impacted performance. Although not formally tested, we observed that large prompts hindered reasoning. Streamlining definitions and instructing Agent4 to respond concisely improved efficiency. These findings align with prior studies showing that excessive or irrelevant context degrades LLM performance [11]. In some cases, we found that sentence order influenced model reasoning, supporting prior work [12].
@@ 4.8 Other Languages
We assess whether the same setup can detect and refactor test smells across different programming languages. Using a single-agent configuration, Phi-4-14B successfully refactored (pass@1) 24/30, 27/30, and 29/30 test smells in Golang, JavaScript, and Python, respectively. These instances contained Assertion Roulette, Duplicate Assert, and Magic Number smells in real-world projects such as kubernetes, react-beautiful-dnd, and django. The model achieved performance comparable to that observed with Java test cases. In the JavaScript setting, it occasionally generated Java code; however, when explicitly instructed to produce JavaScript test code, it responded correctly. JUnit-5 compliance can also be excluded from the prompts. These preliminary results suggest that the proposed approach is promising for detecting and refactoring test smells across a variety of programming languages.
4.9 Detecting Other Types of Test Smells
We evaluated Phi-4-14B on 30 types of test smells [15], using real-world examples—18 in Java and 12 in other languages including Ruby, Python, JavaScript, Smalltalk, and TTCN-3. A single-agent was used, focusing solely on detection, as the catalog does not define refactoring strategies for all smells. Phi-4-14B successfully detected 28 out of 30 test smells (pass@1), providing detailed justifications for each. The two initial failures were due to ambiguous definitions, which, once clarified, enabled the accurate detection of all smells. We also evaluated Phi-4-14B on 45 instances spanning four additional test-smell types in JavaScript. It misclassified only four cases (pass@1) of OverCommented Tests, due to an imprecise definition of when a test contains excessive comments. After clarifying the definition, it correctly detected all test smells.
@@ 4.10 Larger Test Cases
As previously noted, 89% of the test cases in the 11 open-source projects analyzed contain at most 30 lines of code (LOC). To further evaluate Phi-4-14B, which demonstrated the best performance in our study, we selected 30 test smells (five per smell type) exceeding 30 LOC from these 11 open-source projects. We employed the same four-agent setup to detect and refactor these test smells. Although the strategy successfully identified 86% of the test smells, it showed a decline in performance compared to previous results. The effectiveness in accurately refactoring test smells also diminished, with a pass@1 rate of 28%.
@@ 4.11 Threats to Validity
One potential threat to internal validity is data leakage, where test smells analyzed in this study may be part of the foundation models’ training data. To mitigate this, we applied Metamorphic Testing.1 We limit our analysis to test code with a maximum of 30 LOC. While this constraint excludes longer test cases, a substantial number of test cases still fall within this criterion, ensuring a representative evaluation. Another factor is prompt design, which can lead to generic responses instead of targeted explanations. To minimize variability, we used concise prompts across all evaluations, which were carefully reviewed by three authors of this paper. Validating alignment between the model output and test smell definitions is also challenging. To ensure accuracy, three authors independently reviewed selected responses, verifying whether the refactored code adhered to the intended definitions. Construct validity is limited by the small number of examples for each test smell, though they are real cases from GitHub repositories.
@@ 5 Related Work
Aljedaani et al. [3] compiled a comprehensive catalog of 22 test smell detection tools, the majority of which are designed for Java, SmallTalk, C++, and Scala, and 4 refactoring tools. Soares et al. [6] conducted a mixed-methods analysis involving 485 Java projects, exploring the adoption of JUnit-5 features to improve test code quality. Wang et al. [13] propose a tool called PyNose to detect 18 types of test smells in Python. Lambiase et al. [14] introduce DARTS, an IntelliJ plugin that detects and refactors three test smells: Eager Test, General Fixture, and Lack of Cohesion of Test Methods. Pontillo et al. [4] proposed a machine learning-based approach to detect test smells, focusing on four specific types. Lucas et al. [15] investigated the capability of three LLMs to detect test smells across multiple programming languages.
We investigate the effectiveness of agentic workflows for detecting and refactoring test smells in Java test cases. Our approach coordinates up to four-agents to automate these tasks. Results show that Phi-4-14B achieves strong refactoring performance. It runs efficiently on consumer-grade hardware and perform comparably to LLMs. A key advantage of our approach is its extensibility: new smells can be added via natural language definitions. Preliminary experiments also confirm successful application to Python, Golang, and JavaScript, highlighting its cross-language potential. Unlike conventional methods that only detect smells or suggest code changes, our approach supports interactive conversations and explanatory feedback.
@@ 6 Conclusions
We evaluated Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B for the automated detection and refactoring of test smells using agentic workflows with one, two, and four agents. The evaluation covered 150 instances of 5 common test smell types from real-world projects. We assessed each model’s ability to perform automated refactorings. Phi-4-14B, Gemma-2-9B, and Llama-3.2-3B successfully detected nearly all test smell instances. Phi-4-14B delivered the best refactoring performance, reaching a pass@5 of 75.3%, and performed within 5% of proprietary models like o3, Claude-4-Sonnet, and Gemini-2.5-Pro using a single-agent. We submitted 10 pull requests with code refactored by Phi-4-14B. Six were merged, two was rejected, and two remain open – demonstrating the practical utility of our approach in live development environments. Multi-agent workflows outperformed single-agent setups in three out of five test smell types, showing their potential to enhance code quality with minimal developer effort.